If X overbar=65​, S=14​, and n=49​, and assuming that the population is normally​ distributed, construct a 99% confidence interval estimate of the population​ mean, μ.
Answer #1

Ans.

The formula to calculate the 99% confidence interval for population mean is ar{x}pm t_{df=n-1}^{rac{alpha}{2}}left ( rac{s}{sqrt{n}} ight )

t i2 - 2.6822042.6822

Lower limit:

řーぱ(#)- .49 65-2.6822 = 65-5.3644 59.6356

Upper limit:

5+ (#) .49 r48 = 65 + 2.6822 = 65 5.3644 = 70.3644 Vn

So the 99% confidence interval for the population mean is (59.6356,70.3644).